OpenStudy (anonymous):

i have no idea ho to do this prove the trig identity 1-cos(2x)/tan(x) =sin(2x)

OpenStudy (anonymous):

\[1-\cos (2x)/\tan(x)=\sin(2x)\]

terenzreignz (terenzreignz):

What you want to do is to somehow turn the left-hand side of the equation into the right-hand side...

OpenStudy (anonymous):

@terenzreignz stilll dont get it :(

terenzreignz (terenzreignz):

Let's first remove any possible ambiguity \[\Large \frac{1-\cos(2x)}{\tan(x)}=\sin(2x)\] Is it this?

OpenStudy (anonymous):

yes @terenzreignz

terenzreignz (terenzreignz):

Well then, let's see if we can use a little magic... ^_^ We have an established identity for \[\large \cos(2x) =\color{red}?\] What is it?

OpenStudy (anonymous):

cos^2x-sin^2x

OpenStudy (anonymous):

\[\cos^2x-\sin^2x\] @terenzreignz

terenzreignz (terenzreignz):

Okay, let's go with that... so the left-hand side (specifically, the red part) becomes... \[\Large \frac{1-\color{red}{\cos(2x)}}{\tan(x)}=\frac{1-\color{red}{[\cos^2(x)-\sin^2(x)]}}{\tan(x)}\] Okay, catch me so far?

OpenStudy (anonymous):

@terenzreignz a little yes..

terenzreignz (terenzreignz):

Can't be a little, I need you to understand fully before we can proceed... is there something that is unclear to you?

OpenStudy (anonymous):

why did u change sin(2x) to that i thought it would be \[2sinx*cosx\]

terenzreignz (terenzreignz):

It was a cos(2x) remember? Not a sin(2x) We don't care about the sin(2x) yet, our "mission" is to turn the left-hand side of the equation into sin(2x) Which we will... in time... so can we proceed?

OpenStudy (anonymous):

@terenzreignz okok yesss

terenzreignz (terenzreignz):

Okay great, now we simplify... \[\Large \frac{1-\cos^2(x)+\sin^2(x)}{\tan(x)}\] Everything okay so far?

OpenStudy (anonymous):

@terenzreignz what operation did u do? subtraction, division, addition?

terenzreignz (terenzreignz):

I just distributed the minus sign... remember, before that step, it looked like this...\[\Large \frac{1-\color{red}{[\cos^2(x)-\sin^2(x)]}}{\tan(x)}\]

OpenStudy (anonymous):

@terenzreignz okokok! i see

terenzreignz (terenzreignz):

Now this part here... \[\Large \frac{\color{red}{1-\cos^2(x)}+\sin^2(x)}{\tan(x)}\] Does it look familiar to you?

OpenStudy (anonymous):

@terenzreignz yup it doessss!

terenzreignz (terenzreignz):

then it simplifies into...?

terenzreignz (terenzreignz):

Do remember the Pythagorean identity... \[\Large \color{blue}{\sin^2(x)}+\color{red}{\cos^2(x)}= \color{red}1\]

OpenStudy (anonymous):

@terenzreignz NOO :(

terenzreignz (terenzreignz):

then you must... You can't go about proving identities without a thorough knowledge of the basic ones :)

OpenStudy (anonymous):

in my notes it says \[\sin ^2x+\cos ^2x=1\] \[1+\tan ^2x=\sec ^2x\] \[1+\cot ^2x=\csc ^2x\]

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Yeah, those... the first one, is it not what I just wrote? :D

OpenStudy (anonymous):

@terenzreignz ur rightttttt

terenzreignz (terenzreignz):

So, given that first equation, if we subtract \(\cos^2(x)\) from both sides, you get... \[\Large \sin^2(x) =\color{red}{1-\cos^2(x)}\]

OpenStudy (anonymous):

@terenzreignz okok!! yesss

terenzreignz (terenzreignz):

Well then, this part becomes... \[\Large \frac{\color{red}{1-\cos^2(x)}+\sin^2(x)}{\tan(x)}\]

OpenStudy (anonymous):

@terenzreignz OK... so now we got rid of the otherside

terenzreignz (terenzreignz):

No, first, that part in red, it's just equal to \[\Large \frac{\color{red}{\sin^2(x)}+\sin^2(x)}{\tan(x)}\] right?

OpenStudy (anonymous):

@terenzreignz yess righttt

terenzreignz (terenzreignz):

So let's just combine them.. \[\Large \frac{\color{red}{2\sin^2(x)}}{\tan(x)}\]

OpenStudy (anonymous):

@terenzreignz okok!! is there more ways to simplify it?

terenzreignz (terenzreignz):

Of course... \[\large \tan(x) =\frac{\sin(x)}{\cos(x)}\]

terenzreignz (terenzreignz):

So it becomes... \[\Large \frac{2\sin^2(x)}{\frac{\sin(x)}{\cos(x)}}\]

OpenStudy (anonymous):

@terenzreignz ok i get what u did.

terenzreignz (terenzreignz):

So we can bring the cos(x) up... \[\Large = \frac{2\sin^2(x)\cos(x)}{\sin(x)}\]

terenzreignz (terenzreignz):

And finally, cancel out the sin(x) \[\Large = \frac{2\sin^{\cancel2}(x)\cos(x)}{\cancel{\sin(x)}}\]

terenzreignz (terenzreignz):

Leaving you with... \[\Large 2\sin(x)\cos(x)\] Which, obviously, is...?

OpenStudy (anonymous):

1? @terenzreignz

terenzreignz (terenzreignz):

What? NO! \[\Large \sin(2x) = 2\sin(x)\cos(x) \]

OpenStudy (anonymous):

LOL sowwy :(

terenzreignz (terenzreignz):

And that's it... we've successfully turned the left-side into the right side sin(2x)

OpenStudy (anonymous):

@terenzreignz wowww thank u i learned alot!!!!:)

terenzreignz (terenzreignz):

That's all i need to know :)